one mole of a gas has a volume of 0.0223m^3 at a pressure of 101000pa at 0^c. if the molar capacity at constant pressure is 28.5jmol/k.calculate the molar heat capacity at constant volume.​
1
Expert's answer
2020-11-16T07:49:16-0500

Solution

We know relation

"C_p-C_v=R"

Heat capacity at constant volume

"C_v=C_p-R=28.5-8.314"

"C_v=20.186J-mol\/k"